OpenStudy (anonymous):

The volume of a sphere is 5,000pi m3. What is the surface area of the sphere to the nearest square meter?

OpenStudy (anonymous):

surface area of a sphere = \(4\pi r^2\) the volume of a sphere= \(\frac{4}{3} \pi r^3\) so \[5000\pi = \frac{4}{3}\pi r^3\] can you solve for r there ^

OpenStudy (anonymous):

I still dont get it.

OpenStudy (anonymous):

=> r^3 = 5000 x 3 / 4 => r^3 = 3750 Now find the cube root this will give u value of r

OpenStudy (anonymous):

how do you get cube root?

OpenStudy (anonymous):

take LCM of 3750 and make triplets of factors

OpenStudy (anonymous):

so i get 15.5... now what?